[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Fri Jun 16, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Fri Jun 16 18:57:31 UTC 2017


HSU Shift Log for Fri Jun 16 2017 14:57:24 GMT-0400 (EDT)

Shift: H

Forecaster: J Beven

Technical Issues Summary:

ATCF issues: None.

N-AWIPS issues: None.

AWIPS-2/NCP issues: Unable to start any CAVE session on AWIPS-2 on the HSU  
backup/pacific workstation.

Other Techincal issues: None.

Miscellaneous issues: None.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170616/0a61e4bd/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list